OpenStudy (anonymous):

A train leaves New York for Boston, 200 miles away, at 4:00 P.M. and averages 55 mph. Another train leaves Boston for New York on an adjacent set of tracks at 5:00 P.M. and averages 45 mph. At what time will the trains meet? (Round to the nearest minute.)

OpenStudy (anonymous):

d = d1 + d2 d1 = 55 ( t + 1) d2 = 45 t => 55 ( t + 1) + 45 t = 200 miles @arosebailey Can you continue?

OpenStudy (anonymous):

Would t=1.45 ?

OpenStudy (anonymous):

Thus 5 PM + 1hr 27' = 6 hr 27' PM